Find the sum of this infinite geometric series:
100 + 60 + 36 + ...

Answer by stanbon(75887)

Find the sum of this infinite geometric series:
100 + 60 + 36 + ...
---------------
a = 100
r = 0.6
---
S(n) = a/(1-r)
----------------------
S(n) = (100)/(1-0.6) = 100/0.4 = 250
========================================
Cheers,
Stan H.